# Matrix harmonic analysis at high temperature via the Dirichlet process

### v1

@@ 3.2. Approximating measures
For any $\eta\in(0,1]$, we take
$M_{\eta}:=\sup_{u>0}\frac{\log(1+u)}{u^{\eta}}<+\infty$ (3.23)
Then
@@
$C_{\eta,\delta}:=2\min\left\{\frac{M_{1}}{\delta},\frac{M_{\eta}}{\delta^{\eta}}\right\}$ (3.26)
with $M_{\eta}$ given in (3.23) we reclaims (3.20). ∎
@@
[…]
@@ 3.3. Weak convergence in the classical regime
With the notations above, in the regime where $N\to\infty$ and $\beta_{N}\ll N^{-1}$, or equivalently, $c_{N}\to 0^{+}$, we have $\rho_{N}^{(c_{N})}$ converges weakly to $\rho$.
@@
$\int_{\mathbb{R}}\frac{\rho_{N}^{(c_{N})}(\mathrm{d}x)}{(z-x)^{c_{N}}}=\exp\left(-c_{N}\int_{\mathbb{R}}\log(z-s)\rho_{N}(\mathrm{d}s)\right),\quad\text{for all }z\in\mathbb{C}\backslash\mathbb{R}$, (3.35)
@@
$\int_{\mathbb{R}}\frac{\rho_{N}^{(c_{N})}(\mathrm{d}x)}{(z-x)^{c_{N}}}=1-c_{N}\int_{\mathbb{R}}\log(z-s)\rho_{N}^{(c_{N})}(\mathrm{d}s)+\frac{c_{N}^{2}}{2}\int_{\mathbb{R}}\frac{\log(z-s)^{2}}{(z-s)^{\xi}}\rho_{N}^{(c_{N})}(\mathrm{d}s)$ (3.38)
@@
$\begin{split}\int_{\mathbb{R}}\frac{\rho_{N}^{(c_{N})}(\mathrm{d}s)}{z-s}=\int_{\mathbb{R}}\frac{\rho_{N}(\mathrm{d}s)}{z-s}&-\frac{c_{N}}{2}\bigg{(}(2g_{\rho_{N}}(z)-\zeta)g^{\prime}_{\rho_{N}}(z)\exp\left(-\zeta g_{\rho_{N}}(z)\right)\\ &+\int_{\mathbb{R}}\frac{\xi\log(z-s)^{2}-2}{(z-s)^{\xi+1}}\rho_{N}^{(c_{N})}(\mathrm{d}s)\bigg{)}\end{split}$ (3.41)
@@
[…]
$\int_{\mathbb{R}}\frac{\rho_{N}^{(c_{N})}(\mathrm{d}s)}{z-s}=\int_{\mathbb{R}}\frac{\rho_{N}(\mathrm{d}s)}{z-s}+o(1)=\int_{\mathbb{R}}\frac{\rho(\mathrm{d}s)}{z-s}+o(1),\quad\text{ for all }z\in\mathbb{C}\backslash\mathbb{R}$, (3.45)
@@ 3.4. Weak convergence in the high temperature regime
With then notations above, in the regime where $\beta_{N}\ll N^{-1}$ and $N\to\infty$. Equivalently, $c_{N}=c+o(1)$ for some $c>0$, one has $\rho_{N}^{(c_{N})}$ converges weakly to $\rho^{(c)}$ if either of the following assumptions hold: $c>1$ or $\rho$ is compactly supported.
@@
We only discuss the case where $c>1$ or $\rho$ being compactly supported. By Proposition 2.13 we have
@@
One aims at showing those two quantities are close as $N\to\infty$.
For convenience, let us denote
@@
$\left|\int_{\mathbb{R}}\rho_{N}^{(c_{N})}(x)e^{itx}-\int_{\mathbb{R}}\rho^{(c)}(x)e^{itx}\right|$ $\leq\left|\frac{\Gamma(c_{N})}{(it)^{c_{N}-1}}\frac{1}{2\pi i}\right|\cdot\left|F_{N}(t)-F(t)\right|+\left|\frac{\Gamma(c_{N})}{(it)^{c_{N}-1}}-\frac{\Gamma(c)}{(it)^{c-1}}\right|\cdot\left|\frac{1}{2\pi i}F(t)\right|$ (3.51)
The second term can be bounded by
@@
$\leq\left|\frac{\Gamma(c_{N})}{\Gamma(c)}(it)^{c-c_{N}}-1\right|\to 0.$ (3.53)
Also since $\left|\frac{\Gamma(c_{N})}{(it)^{c_{N}-1}}\frac{1}{2\pi i}\right|$ is bounded, it remains to show the difference between $F$ and $F_{N}$ is small. For this purpose, using the identity
@@
$\left|F_{N}(t)-F(t)\right|$ $=\left|\int_{\mathbb{R}-i\delta\mathrm{sgn}(t)}\mathrm{d}z\,\exp\left(itz-cg_{\rho}(z)\right)\left(\exp\left(-c_{N}g_{\rho_{N}-\rho}(z)-(c_{N}-c)g_{\rho}(z)\right)-1\right)\right|.$ (3.55)
@@
Lemma 3.9.
With the above notaions,
$\left|\exp\left(-c_{N}g_{\rho_{N}-\rho}(z)-(c_{N}-c)g_{\rho}(z)\right)-1\right|\leq\left(b_{N}+\frac{|c_{N}-c|}{2}\log(x^{2}+\delta^{2})\right)e^{b_{N}}(x^{2}+\delta^{2})^{|c_{N}-c|/2}$ (3.56)
[…]
@@
$|g_{\rho_{N}-\rho}(z)|\leq C_{\eta,\delta}d_{W_{\eta}}(\rho_{N},\rho),\quad|g_{\rho}(z)|\leq\log|z|+K_{\delta,\rho}.$ (3.59)
@@
$\begin{split}&\left|\exp\left(-c_{N}g_{\rho_{N}-\rho}(z)-(c_{N}-c)g_{\rho}(z)\right)-1\right|\\ &\quad\leq\left(c_{N}C_{\eta,\delta}d_{W_{\eta}}(\rho_{N},\rho)+|c_{N}-c|(\log|z|+K_{\delta,\rho})\right)e^{c_{N}C_{\eta,\delta}d_{W_{\eta}}(\rho_{N},\rho)+|c_{N}-c|K_{\delta,\rho}}\,|z|^{|c_{N}-c|}.\end{split}$ (3.60)
@@
$b_{N}:=c_{N}C_{\eta,\delta}d_{W_{\eta}}(\rho_{N},\rho)+|c_{N}-c|K_{\delta,\rho}.$ (3.61)
Note that $b_{N}$ is independent of $z$. With $z=x-i\delta\mathrm{sgn}(t)$, we have $|z|=(x^{2}+\delta^{2})^{1/2}$ and reclaim the result. ∎
Lemma 3.10.
For all $c>0,\delta\geq 1$, $\varepsilon\in(0,1-\frac{1}{c})$, $x\in\mathbb{R}$ and a probability measure $\rho$ on $\mathbb{R}$, there exists a constant $T_{\varepsilon,\rho}$ such that for all $T>T_{\varepsilon,\rho}$, there are constants $C_{T,\delta}>0$ and $M_{T,\rho}\in(\varepsilon+\frac{1}{c},1)$ given in (3.71) and (3.66) such that
$\begin{split}&\exp\left(-c\int_{\mathbb{R}}\frac{1}{2}\log(\delta^{2}+(x-s)^{2})\rho(\mathrm{d}s)\right)\leq(C_{T,\delta})^{cM_{T,\rho}}(x^{2}+\delta^{2})^{-cM_{T,\rho}/2}.\end{split}$ (3.62)
@@
For $T>0$, we denote the function
$f_{T}(x)=\begin{cases}\frac{1}{2}\log(\delta^{2}+(x+T)^{2}),&x<-T,\\ \log\delta,&-T\leq x\leq T,\\ \frac{1}{2}\log(\delta^{2}+(x-T)^{2}),&x>T.\end{cases}$ (3.63)
It can be checked that for all $\delta\geq 1$ and $s\in(-T,T)$, one has
$0\leq f_{T}(x)\leq\frac{1}{2}\log(\delta^{2}+(x-s)^{2}),\quad x\in\mathbb{R}.$ (3.64)
Therefore,
$\int_{\mathbb{R}}\frac{1}{2}\log(\delta^{2}+(x-s)^{2})\rho(\mathrm{d}s)$ $\geq\int_{-T}^{T}\frac{1}{2}\log(\delta^{2}+(x-s)^{2})\rho(\mathrm{d}s)\geq f_{T}(x)\int_{-T}^{T}\rho(\mathrm{d}s)\geq 0$ (3.65)
By the Markov inequality one can furthermore give an lower bound
$\int_{-T}^{T}\rho(\mathrm{d}s)\geq 1-\frac{\int_{\mathbb{R}}\log(1+s^{2})\rho(\mathrm{d}s)}{\log(1+T^{2})}=:M_{T,\rho}.$ (3.66)
For a given $\rho$ with $\int_{\mathbb{R}}\log(1+s^{2})\rho(\mathrm{d}s)<+\infty$, one can take a $T$ large enough ($T\geq T_{\varepsilon,\rho}$ for some $T_{\varepsilon,\rho}$) so that
$\frac{1}{c}+\varepsilon<M_{T,\rho}<1,\quad\text{(recall that $c>1$)},$ (3.67)
for all $\varepsilon\in(0,1-\frac{1}{c})$. Thus one has
@@
$\exp\left(-c\int_{\mathbb{R}}\frac{1}{2}\log(\delta^{2}+(x-s)^{2})\rho(\mathrm{d}s)\right)\leq\left(\exp(-f_{T}(x))\right)^{cM_{T,\rho}}.$ (3.68)
Now we will give the following bound for $\exp(-f_{T}(x))$
$\exp(-f_{T}(x))\leq C_{T}(x^{2}+\delta^{2})^{-1/2}$ (3.69)
for some $C_{T}>0$. The above inequality can be simplfied to
$C_{T}^{2}\geq g(x):=\begin{cases}\displaystyle\frac{x^{2}+\delta^{2}}{(x+T)^{2}+\delta^{2}},&x<-T;\vskip 5.0pt\\ \displaystyle\frac{x^{2}+\delta^{2}}{\delta^{2}},&-T\leq x\leq T;\vskip 5.0pt\\ \displaystyle\frac{x^{2}+\delta^{2}}{(x-T)^{2}+\delta^{2}},&x>T.\end{cases}$ (3.70)
The maximum values of $g$ in each of the three domains can be computed explicitly. Thus,
@@
$C_{T}\geq\max\left(1+\frac{T^{2}}{\delta^{2}},\frac{1}{2}\left(1+\sqrt{1+\frac{4\delta^{2}}{T^{2}}}\right)\right)^{1/2}=:C_{T,\delta},$ (3.71)
Substituting these two bounds back gives
$\left(\exp(-f_{T}(x))\right)^{cM_{T,\rho}}\leq(C_{T,\delta})^{cM_{T,\rho}}(x^{2}+\delta^{2})^{-cM_{T,\rho}/2}$ (3.72)
which is the desired result. ∎

### v3

@@ 3.2. Approximating measures
For any $\eta\in(0,1]$, we take an upper bound
$\sup_{u>0}\frac{\log(1+u)}{u^{\eta}}\leq\frac{e^{\eta-1}}{\eta}=:M_{\eta}.$ (3.23)
To verify this, in the case $|u|\leq 1$ it is
$\frac{\log(1+u)}{u^{\eta}}\leq u^{1-\eta}\leq 1,$ (3.24)
and for $|u|>1$ one has
$\frac{\log(1+u)}{u^{\eta}}\leq\frac{\log(eu)}{u^{\eta}}=\frac{1+\log u}{u^{\eta}}\leq\frac{e^{\eta-1}}{\eta}.$ (3.25)
Also since $\frac{e^{\eta-1}}{\eta}\geq 1$, combining the previous two bounds we claim (3.23). Then one has
@@
$C_{\eta,\delta}:=2\max\left\{\frac{M_{1}}{\delta},\frac{M_{\eta}}{\delta^{\eta}}\right\}$ (3.28)
with $M_{\eta}$ given in (3.23) we reclaims (3.20). For future use, we remark that when $\delta=1$,
$C_{\eta,1}=2\max\left\{1,\frac{e^{\eta-1}}{\eta}\right\}\leq\frac{e^{\eta-1}}{\eta}.$ (3.29)
@@
For a given $z\in\mathbb{C}\setminus\mathbb{R}$, the following inequality holds
@@ 3.3. Weak convergence in the classical regime
With the notations above, in the regime where $N\to+\infty$ and $\beta_{N}\ll N^{-1}$, or equivalently, $c_{N}\to 0^{+}$, we have $\rho_{N}^{(c_{N})}$ converges weakly to $\rho$.
@@
$\int_{\mathbb{R}}\frac{\rho_{N}^{(c_{N})}(\mathrm{d}x)}{(z-x)^{c_{N}}}=\exp\left(-c_{N}\int_{\mathbb{R}}\log(z-s)\rho_{N}(\mathrm{d}s)\right),\quad\text{for all }z\in\mathbb{C}\setminus\mathbb{R},$ (3.38)
@@
$\int_{\mathbb{R}}\frac{\rho_{N}^{(c_{N})}(\mathrm{d}x)}{(z-x)^{c_{N}}}=1-c_{N}\int_{\mathbb{R}}\log(z-s)\rho_{N}^{(c_{N})}(\mathrm{d}s)+\frac{c_{N}^{2}}{2}\int_{\mathbb{R}}\frac{\log(z-s)^{2}}{(z-s)^{\xi}}\rho_{N}^{(c_{N})}(\mathrm{d}s).$ (3.41)
@@
$\begin{split}\int_{\mathbb{R}}\frac{\rho_{N}^{(c_{N})}(\mathrm{d}s)}{z-s}=\int_{\mathbb{R}}\frac{\rho_{N}(\mathrm{d}s)}{z-s}&-\frac{c_{N}}{2}\bigg((2g_{\rho_{N}}(z)-\zeta)g^{\prime}_{\rho_{N}}(z)\exp\left(-\zeta g_{\rho_{N}}(z)\right)\\ &+\int_{\mathbb{R}}\frac{\xi\log(z-s)^{2}-2}{(z-s)^{\xi+1}}\rho_{N}^{(c_{N})}(\mathrm{d}s)\bigg).\end{split}$ (3.44)
@@
in the $N\to+\infty$ (i.e. $c_{N}\to 0^{+}$) limit, pointwise for $z\in\mathbb{C}\setminus\mathbb{R}$.
Now we make use of the convergence theorem for Stieltjes transform (see e.g. [4, Theorem 2.4.4]): a sequence of probability measure weakly converges to another probability measure, if and only if their Stieltjes transforms converge pointwise for all $z\in\mathbb{C}\setminus\mathbb{R}$. Noting that
$\int_{\mathbb{R}}\frac{\rho_{N}^{(c_{N})}(\mathrm{d}s)}{z-s}=\int_{\mathbb{R}}\frac{\rho_{N}(\mathrm{d}s)}{z-s}+o(1)=\int_{\mathbb{R}}\frac{\rho(\mathrm{d}s)}{z-s}+o(1),\quad\text{ for all }z\in\mathbb{C}\setminus\mathbb{R},$ (3.48)
@@ 3.4. Weak convergence in the high temperature regime
With then notations above, in the regime where $\beta_{N}\ll N^{-1}$ and $N\to+\infty$. Equivalently, $c_{N}=c+o(1)$ for some $c>0$, one has $\rho_{N}^{(c_{N})}$ converges weakly to $\rho^{(c)}$ for all $c>1$.
@@
We only discuss the case where $c>1$. By Proposition 2.9 we have
@@
One aims at showing those two quantities are close as $N\to+\infty$. For convenience, let us denote
@@
$\left|\int_{\mathbb{R}}\rho_{N}^{(c_{N})}(x)e^{itx}-\int_{\mathbb{R}}\rho^{(c)}(x)e^{itx}\right|$ $\leq\underbrace{\left|\frac{\Gamma(c_{N})}{(it)^{c_{N}-1}}\frac{1}{2\pi i}\right|}_{I_{1}}\cdot\underbrace{\left|F_{N}(t)-F(t)\right|}_{I_{2}}+\underbrace{\left|\frac{\Gamma(c_{N})}{(it)^{c_{N}-1}}-\frac{\Gamma(c)}{(it)^{c-1}}\right|\cdot\left|\frac{1}{2\pi i}F(t)\right|}_{I_{3}}.$ (3.54)
To give an upper bound of this quantity, it requires to give bounds to each terms $I_{1},I_{2}$ and $I_{3}$.
One firstly states the following lemma for bounds for $I_{1}$ and $I_{3}$.
Lemma 3.9 (Bounds for $I_{1}$ and $I_{3}$).
If $|c_{N}-c|<\varepsilon$ for some $\varepsilon>0$, then
$I_{1}=\left|\frac{\Gamma(c_{N})}{(it)^{c_{N}-1}}\frac{1}{2\pi i}\right|\leq\frac{\max\{\Gamma(c-\varepsilon),\Gamma(c+\varepsilon)\}|t|^{|c+\varepsilon-1|}}{2\pi}$ (3.55)
and
$I_{3}=\left|\frac{\Gamma(c_{N})}{(it)^{c_{N}-1}}-\frac{\Gamma(c)}{(it)^{c-1}}\right|\cdot\left|\frac{1}{2\pi i}F(t)\right|\leq|c_{N}-c|B,$ (3.56)
where
$B:=\left(\sup_{|w-c|\leq\epsilon}|\psi(w)|+|\log|t||+\frac{\pi}{2}\right)e^{\varepsilon\left(\sup_{|w-c|\leq\epsilon}|\psi(w)|+|\log|t||+\frac{\pi}{2}\right)}.$ (3.57)
Proof.
For $I_{1}$, the bound can be observed by notating that the Gamma function is convex on $\mathbb{R}^{+}$ and so $\Gamma(c_{n})\leq\max\{\Gamma(c-\varepsilon),\Gamma(c+\varepsilon)\}$, and that
$|t^{-(c_{N}-1)}|\leq\exp|(c_{N}-1)\log|t||\leq\exp|(c+\varepsilon-1)\log|t||.$ (3.58)
For $I_{3}$, we have
@@
$\leq\left|\frac{\Gamma(c_{N})}{\Gamma(c)}(it)^{c-c_{N}}-1\right|.$ (3.60)
It remains to give an upper bound of the right side. Taking
$z:=\log\Gamma(c_{N})-\log\Gamma(c)+(c-c_{N})\log(it),$ (3.61)
we can use the notion of digamma function $\psi(u)=(\log\Gamma(u))^{\prime}$ to obtain
$z=(c-c_{N})\left(\int_{0}^{1}\psi(c+u(c-c_{N}))\mathrm{d}u-\log(it)\right).$ (3.62)
Since $\psi$ is analytic near $c>1$, the quantity $\sup_{|w-c|\leq\epsilon}|\psi(w)|<+\infty$ is bounded and one has
$|z|\leq|c_{N}-c|\left(\sup_{|w-c|\leq\epsilon}|\psi(w)|+|\log|t||+\frac{\pi}{2}\right).$ (3.63)
By using the inequalities $|e^{z}-1|\leq|z|e^{|z|}$ and
$\left|\frac{1}{2\pi i}F(t)\right|=\left|\int_{\mathbb{R}}e^{itx}\rho^{(c)}(\mathrm{d}x)\right|\leq 1$ (3.64)
we have
$\left|\frac{\Gamma(c_{N})}{\Gamma(c)}(it)^{c-c_{N}}-1\right|\leq|c_{N}-c|\left(\sup_{|w-c|\leq\epsilon}|\psi(w)|+|\log|t||+\frac{\pi}{2}\right)e^{\varepsilon\left(\sup_{|w-c|\leq\epsilon}|\psi(w)|+|\log|t||+\frac{\pi}{2}\right)}$ (3.65)
which finishes the proof. ∎
It remains to show $I_{2}$, the difference between $F$ and $F_{N}$, is small. For this purpose, using the identity
@@
$\left|F_{N}(t)-F(t)\right|$ $=\left|\int_{\mathbb{R}-i\delta\mathrm{sgn}(t)}\mathrm{d}z\,\underbrace{\exp\left(itz-cg_{\rho}(z)\right)}_{I_{4}}\underbrace{\left(\exp\left(-c_{N}g_{\rho_{N}-\rho}(z)-(c_{N}-c)g_{\rho}(z)\right)-1\right)}_{I_{5}}\right|.$ (3.67)
@@
Lemma 3.10 (Bound for $I_{5}$).
With the above notations, for any $\varepsilon>0$, one has
$\left|\exp\left(-c_{N}g_{\rho_{N}-\rho}(z)-(c_{N}-c)g_{\rho}(z)\right)-1\right|\leq\left(\frac{b_{N}}{\delta^{\varepsilon}}+\frac{|c_{N}-c|}{2e\varepsilon}\right)e^{b_{N}}(x^{2}+\delta^{2})^{|c_{N}-c|/2+\varepsilon}$ (3.68)
with $z=x-i\delta\mathrm{sgn}(t)$, $x\in\mathbb{R}$, $\delta>0$ and $t\in\mathbb{R}\setminus\{0\}$, where the constant $b_{N}$ is given by (3.73).
@@
$|g_{\rho_{N}-\rho}(z)|\leq C_{\eta,\delta}\mathrm{d}_{W_{\eta}}(\rho_{N},\rho),\quad|g_{\rho}(z)|\leq\log|z|+K_{\delta,\rho}.$ (3.71)
@@
$\begin{split}&\left|\exp\left(-c_{N}g_{\rho_{N}-\rho}(z)-(c_{N}-c)g_{\rho}(z)\right)-1\right|\\ &\quad\leq\left(c_{N}C_{\eta,\delta}\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)+|c_{N}-c|(\log|z|+K_{\delta,\rho})\right)e^{c_{N}C_{\eta,\delta}\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)+|c_{N}-c|K_{\delta,\rho}}\,|z|^{|c_{N}-c|}.\end{split}$ (3.72)
@@
$b_{N}:=c_{N}C_{\eta,\delta}\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)+|c_{N}-c|K_{\delta,\rho}.$ (3.73)
Note that $b_{N}$ is independent of $z$. With $z=x-i\delta\mathrm{sgn}(t)$, we have $|z|=(x^{2}+\delta^{2})^{1/2}$ and so
$\left|\exp\left(-c_{N}g_{\rho_{N}-\rho}(z)-(c_{N}-c)g_{\rho}(z)\right)-1\right|\leq\left(b_{N}+\frac{|c_{N}-c|}{2}\log(x^{2}+\delta^{2})\right)e^{b_{N}}(x^{2}+\delta^{2})^{|c_{N}-c|/2}.$ (3.74)
To simplify the previous inequality, one takes the inequalities
$b_{N}e^{b_{N}}(x^{2}+\delta^{2})^{|c_{N}-c|/2}\leq\frac{b_{N}e^{b_{N}}}{\delta^{\varepsilon}}(x^{2}+\delta^{2})^{|c_{N}-c|/2+\varepsilon},$ (3.75)
$\frac{|c_{N}-c|}{2}e^{b_{N}}\log(x^{2}+\delta^{2})\leq\frac{|c_{N}-c|}{2e\varepsilon}e^{b_{N}}(x^{2}+\delta^{2})^{\varepsilon},$ (3.76)
which lead to the final result. ∎
Proposition 3.11 (Bound for $I_{4}$).
Fix $\delta>0$ and $c>0$. Then for every $M>0$ there exist a constant $A$ such that for all $x\in\mathbb{R}$
$\exp\left(-\frac{c}{2}\int_{\mathbb{R}}\log((x-s)^{2}+\delta^{2})\,\rho(\mathrm{d}s)\right)\leq e^{-\frac{cA}{2}}(x^{2}+\delta^{2})^{-\tfrac{c\rho([-M,M])}{2}}.$ (3.77)
where
$A:=\rho([-M,M])\log\min\left\{\frac{1}{4},\frac{\delta^{2}}{\delta^{2}+4M^{2}}\right\}+\rho(\mathbb{R}\setminus[-M,M])\log\delta^{2}.$ (3.78)
@@
Let us decompose
$J(x):=\int_{\mathbb{R}}\log((x-s)^{2}+\delta^{2})\,\rho(\mathrm{d}s)=J_{\mathrm{in}}(x)+J_{\mathrm{out}}(x),$ (3.79)
where
$J_{\mathrm{in}}(x):=\int_{|s|\leq M}\log((x-s)^{2}+\delta^{2})\,\rho(\mathrm{d}s),\qquad J_{\mathrm{out}}(x):=\int_{|s|>M}\log((x-s)^{2}+\delta^{2})\,\rho(\mathrm{d}s).$ (3.80)
Step 1 is to show a lower bound for $J_{\mathrm{in}}(x)$. For each $|s|\leq M$ and every $x\in\mathbb{R}$, we will check that
$(x-s)^{2}+\delta^{2}\geq\min\left\{\frac{1}{4},\frac{\delta^{2}}{\delta^{2}+4M^{2}}\right\}(x^{2}+\delta^{2}).$ (3.81)
Let us show (3.81) first. In the case $|x|<2M$, one has
$(x-s)^{2}+\delta^{2}\geq\delta^{2}\geq\frac{\delta^{2}}{\delta^{2}+4M^{2}}(x^{2}+\delta^{2}).$ (3.82)
In the case $|x|\geq 2M$, one has $|x|-M\geq\frac{x}{2}$ and so a triangualar inequality gives
$(x-s)^{2}\geq||x|-M|^{2}\geq\frac{x^{2}}{4}.$ (3.83)
This implies
$(x-s)^{2}+\delta^{2}\geq\delta^{2}+\frac{x^{2}}{4}\geq\frac{1}{4}(\delta^{2}+x^{2}).$ (3.84)
Combining the two inequalities (3.82) and (3.84) we find the constant $\min\left\{\frac{1}{4},\frac{\delta^{2}}{\delta^{2}+4M^{2}}\right\}$. Taking logarithms and integrating over $\rho$ for (3.81) gives
$J_{\mathrm{in}}(x)\geq\rho([-M,M])\log(x^{2}+\delta^{2})+\rho([-M,M])\log\min\left\{\frac{1}{4},\frac{\delta^{2}}{\delta^{2}+4M^{2}}\right\}.$ (3.85)
Next we will show a lower bound for $J_{\mathrm{out}}(x)$. Trivially, $(x-s)^{2}+\delta^{2}\geq\delta^{2}$ for all $s,x$, hence
$J_{\mathrm{out}}(x)\geq\rho(\mathbb{R}\setminus[-M,M])\log\delta^{2}.$ (3.86)
Combining (3.85) and (3.86) gives
$J(x)\geq\rho([-M,M])\log(x^{2}+\delta^{2})+\rho([-M,M])\log\min\left\{\frac{1}{4},\frac{\delta^{2}}{\delta^{2}+4M^{2}}\right\}+\rho(\mathbb{R}\setminus[-M,M])\log\delta^{2}.$ (3.87)
Exponentiating this gives the desired result. ∎
Lemma 3.12 (Bound for $I_{2}$).
We take $\delta=1$ and $c>1$. For any $\varepsilon\in(0,\frac{c-1}{5})$, $M>0$ with $\rho([-M,M])\in(\frac{1+5\varepsilon}{c},1]$ and $N>N_{0}$ (such that $|c_{N}-c|<\varepsilon$), one has
$\left|F_{N}(t)-F(t)\right|\leq 2\sqrt{\frac{\pi}{\varepsilon}}e^{\frac{c}{2}\log(1+4M^{2})+|t|}\left(b_{N}+\frac{|c_{N}-c|}{2e\varepsilon}\right)e^{b_{N}}$ (3.88)
where
$b_{N}\leq\frac{(c+\varepsilon)e^{\eta-1}}{\eta}\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)+|c_{N}-c|\left(\frac{1}{2}\int_{\mathbb{R}}\log(1+s^{2})\rho(\mathrm{d}s)+4\right),\quad\eta\in(0,1).$ (3.89)
Proof.
Lemma (3.11) shows that for $z\in\mathbb{R}-i\delta\mathrm{sgn}(t)$
$\left|\exp\left(itz-cg_{\rho}(z)\right)\right|\leq e^{-\frac{cA}{2}}(x^{2}+\delta^{2})^{-c\rho([-M,M])/2}e^{|t|\delta}.$ (3.90)
Substituting the inequality in Lemma 3.10 and (3.90) into (3.67) gives
@@
$\begin{split}\left|F_{N}(t)-F(t)\right|&\leq e^{-\frac{cA}{2}+|t|\delta}\left(\frac{b_{N}}{\delta^{\varepsilon}}+\frac{|c_{N}-c|}{2e\varepsilon}\right)e^{b_{N}}\int_{\mathbb{R}}\mathrm{d}x\,(x^{2}+\delta^{2})^{-c\rho([-M,M])/2+|c_{N}-c|/2+\varepsilon}.\end{split}$ (3.91)
As $c>1$, one takes $\varepsilon\in(0,\frac{c-1}{5})$ and $M>0$ with $\rho([-M,M])\in(\frac{1+5\varepsilon}{c},1]$. Then there exists $N_{0}$ such that for all $N>N_{0}$ one has $|c_{N}-c|<\varepsilon$. Then the inequality
$\alpha_{N}:=c\rho([-M,M])/2-|c_{N}-c|/2-\varepsilon\in\left(\frac{1}{2}+\varepsilon,\frac{c}{2}\right).$ (3.92)
holds. So the previous integral is absolute integrable and can be computed as
$\begin{split}\int_{\mathbb{R}}\mathrm{d}x\,(x^{2}+\delta^{2})^{-\alpha}=\frac{\sqrt{\pi}\Gamma(\alpha_{N}-\frac{1}{2})}{\Gamma(\alpha_{N})}\delta^{1-2\alpha_{N}}.\end{split}$ (3.93)
Note the two inequalities simplify the bound
@@
$\left|\delta^{1-2\alpha_{N}}\right|\leq 1+\delta^{1-c},\qquad\left|\frac{\sqrt{\pi}\Gamma(\alpha_{N}-\frac{1}{2})}{\Gamma(\alpha_{N})}\right|\leq\sqrt{\frac{\pi}{\alpha_{N}-1/2}}\leq\sqrt{\frac{\pi}{\varepsilon}}.$ (3.94)
We thus have
$\left|F_{N}(t)-F(t)\right|\leq\sqrt{\frac{\pi}{\varepsilon}}(1+\delta^{1-c})e^{-\frac{cA}{2}+|t|\delta}\left(\frac{b_{N}}{\delta^{\varepsilon}}+\frac{|c_{N}-c|}{2e\varepsilon}\right)e^{b_{N}}.$ (3.95)
To further simplify this bound, we take $\delta=1$ (by the arbitrariness of $\delta$) and $M>\delta$. Then
$A\geq-\log(1+4M^{2})$ (3.96)
and one thus has
$\left|F_{N}(t)-F(t)\right|\leq 2\sqrt{\frac{\pi}{\varepsilon}}e^{\frac{c}{2}\log(1+4M^{2})+|t|}\left(b_{N}+\frac{|c_{N}-c|}{2e\varepsilon}\right)e^{b_{N}},$ (3.97)
where
$b_{N}\leq\frac{(c+\varepsilon)e^{\eta-1}}{\eta}\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)+|c_{N}-c|\left(\frac{1}{2}\int_{\mathbb{R}}\log(1+s^{2})\rho(\mathrm{d}s)+4\right),\quad\eta\in(0,1).$ (3.98)
(we use the inequality $\frac{1}{2}\log 2+\pi<4$ only for convenience). ∎